OpenStudy (anonymous):

Evaluate: 1/8 + 2/3 x 4 would that be 2 19/24

OpenStudy (anonymous):

i got 67/24

OpenStudy (anonymous):

easiest thing to do is change them all in to something over 24: 3/24 +16/24 +96/24=(3+16+96)/24=115/24

sam (.sam.):

@Brooke3001 yes you're correct

OpenStudy (anonymous):

wait how?

OpenStudy (anonymous):

@sheena101 well you are completly right with the 67/24 but then you have to turn it into a mixed number by seeing how many times 24 goes into 67 which would equal 2 19/24

sam (.sam.):